OpenStudy (anonymous):

Mr.bob will select 2 of the 12 students in the math club to be members of the math team. how many different 2 student teams can be chosen from the 12 in the club???

OpenStudy (anonymous):

its 12C2

OpenStudy (anonymous):

\[\dbinom{12}{2}=\frac{12\times 11}{2}=6\times 11=66\]

OpenStudy (anonymous):

satellite,wat happened to my question ? :(